# Bertrand’s and Rodriguez Villegas’ Conjecture for real quartic Galois extensions of the rationals

### v1

@@ Abstract.
The conjecture due to Bertrand and Rodriguez Villegas asserts that the 1-norm of the nonzero element in an exterior power of the units of a number field has a certain lower bound. For the exterior square case of totally real quartic extensions of the rationals, Costa and Friedman gave a lower bound of 0.802. We prove that the bound can be improved to 1.203 when the extension is further assumed to be Galois.
@@
Bertrand–Rodriguez Villegas conjecture, units
@@ 1. Introduction
While Dirichlet’s theorem determines the rank of the unit group in a given algebraic number field in terms of its Archimedean places, less is known about its geometry when viewed as a lattice in a Euclidean space via the associated logarithm embedding. The resulting lattice and its exterior powers are conjecturally subject to certain numerical constraints [1, 2]. We begin by introducing notation to clarify the nature of these numerical constraints.
@@
to get an orthonormal basis  $\{\delta^{I}\}_{I\in\mathcal{A}_{L}^{[j]}}$ of  $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$. For $w=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}\delta^{I}\in\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$, define its 1-norm as
$\lVert w\rVert_{1}:=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}\lvert c_{I}\rvert.$
Our aim in this paper is to investigate the lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of  $\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. For a positive integer $n$, let $L_{n}$ denote the set of subfields of $\mathbb{R}$ that are Galois extension of $\mathbb{Q}$ of degree $n$. For such fields, $\operatorname{rank}({\mathcal{O}_{L}^{*}})=n-1$ by Dirichlet’s unit theorem. For $11\leq j\leq n-$ 1, define a real number $A_{n,j}$ as
@@
This constant is directly related to the real Galois extension case of the Bertrand’s and Rodriguez Villegas’ conjecture. The conjecture stated in [2] is as follows.
@@
There exists two absolute constants
@@
$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802,$
as explained in § 2. To the best of the author’s knowledge, this was the largest known lower bound for $A_{4,2}$.
@@
$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203$
@@
$A_{4,2}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203.$
As demonstrated in [2], Bertrand–Rodriguez Villegas conjecture is equivalent to Lehmer’s conjecture [6] when $j=1$, and to Zimmert’s theorem on regulators [8] when $j=\operatorname{rank}({\mathcal{O}_{L}^{*}})$. Hence the case of  $n=4,j=2$ is neither covered by Lehmer’s conjecture nor Zimmert’s theorem. While our result only covers this specific case, we obtained a better lower bound, which is $1.5$ times the previous known lower bound. In fact, for $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $w=\pm\operatorname{LOG}(\frac{1+\sqrt{5}}{2})\wedge\operatorname{LOG}(1+\sqrt{2})$, we have  $\lVert w\rVert_{1}=8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})$ and thus
$A_{4,2}\leq 8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})\approx 3.3930.$
Hence our lower bound is no less than a third of an upper bound of $A_{4,2}$.
The proof of the theorem is done separately for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is Klein four-group, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is cyclic. For the Klein four-group case, we define a subgroup $E$ generated by the units of the quadratic subfields of $L$. Using the Galois module structure of $E$, we give a lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of $\bigwedge^{2}\operatorname{LOG}(E)$ and extend this result to ${\mathcal{O}_{L}^{*}}$. For the cyclic group case, we use the structure of  ${\mathcal{O}_{L}^{*}}$ given by Hasse [5] to obtain a basis of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. We then use the result of Pohst [7] to compute the lower bound.
In § 2 we derive the best lower bound of  $A_{4,2}$ known so far, from the work of Costa and Friedman [3]. In the remaining sections, we prove the main theorem to give the better lower bound for $A_{4,2}$. The case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group is covered in § 3, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group is covered in § 4.
@@ Acknowledgement
The work of S.S. was supported by College of Natural Sciences Undergraduate Internship Program from Seoul National University. D.K. is supported by the National Research Foundation of Korea (NRF)111the grants No. RS-2023-00301976 and No. 2020R1C1C1A0100681913.
@@ 2. The lower bound due to Costa and Friedman
In this section, we derive the largest lower bound of  $A_{4,2}$ known so far, to the best of the author’s knowledge. Using Pohst’s result [7], Costa and Friedman [3] showed that for independent elements $\epsilon_{1},...,\epsilon_{j}$ of ${\mathcal{O}_{L}^{*}}$, the inequality
$\lVert\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{j})\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
holds for $1\leq j<n$, where $n=[L:\mathbb{Q}]$ and  $\gamma_{j}$ is Hermite’s constant in dimension $j$. If $j=n-2=\operatorname{rank}({\mathcal{O}_{L}^{*}})-1$, every element of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ can be written as  $d\cdot\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{n-2})$ where $d\in\mathbb{Z}$ and $\operatorname{LOG}(\epsilon_{1}),...,\operatorname{LOG}(\epsilon_{n-1})$ are the basis of $\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, as shown in the Lemma 28 of [2]. Therefore, for any nonzero element  $w$ of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, the inequality
$\lVert w\rVert_{1}\geq\lVert w\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
@@
$A_{n,n-2}>\left(\frac{n}{\gamma_{n-2}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{\frac{n-2}{2}}$
@@
$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802.$
@@ 3. The Klein four-group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Let $\mathbb{Q}(\sqrt{d_{1}}),\mathbb{Q}(\sqrt{d_{2}}),\mathbb{Q}(\sqrt{d_{3}})$ be three quadratic subfields of $L$, where $d_{1},d_{2},d_{3}$ are square-free integers. Then we can write $L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$. For $i=1,2,3$, let $\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))=\{1,\sigma_{i}\}$. See Figure 1. Let  $u_{i}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{d_{i}})$. Without loss of generality, let $1<u_{1}<u_{2}<u_{3}$. If $i$ and $j$ are distinct integers in $\{1,2,3\}$, since $\sqrt{d_{j}}\not\in\mathbb{Q}(\sqrt{d_{i}})$$\sigma_{i}\in\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))$ sends $\sqrt{d_{j}}$ to its conjugate $-\sqrt{d_{j}}$. Thus
@@
$\mathbb{Q}$$\mathbb{Q}(\sqrt{d_{2}})$$\mathbb{Q}(\sqrt{d_{3}})$$\mathbb{Q}(\sqrt{d_{1}})$$L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$$\{id,\sigma_{1}\}$$\{id,\sigma_{2}\}$$\{id,\sigma_{3}\}$$\{id,\sigma_{1},\sigma_{2},\sigma_{3}\}$$\{id\}$
Figure 1. Subfields of non-cyclic extension $L$ and corresponding Galois groups
@@
Proof.
Assume that
$u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}=1$
for some $m_{1},m_{2},m_{3}\in\mathbb{Z}$. Then we have
$\sigma_{1}(u_{1})^{m_{1}}\sigma_{1}(u_{2})^{m_{2}}\sigma_{1}(u_{3})^{m_{3}}=1$
and thus
$(u_{1}\sigma_{1}(u_{1}))^{m_{1}}(u_{2}\sigma_{1}(u_{2}))^{m_{2}}(u_{3}\sigma_{1}(u_{3}))^{m_{3}}=1.$
By (3.1), $u_{2}\sigma_{1}(u_{2})=\pm 1$ and $u_{3}\sigma_{1}(u_{3})=\pm 1$. Therefore
$u_{1}^{2m_{1}}=\pm 1$
and thus $m_{1}=0$. Similarly, $m_{2}=m_{3}=0$. Therefore, $u_{1},u_{2},u_{3}$ are multiplicatively independent. ∎
Let $E=\{u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}|m_{1},m_{2},m_{3}\in\mathbb{Z}\}$, then we have the following lemma.
@@
Proof.
From  $u\in{\mathcal{O}_{L}^{*}}$, we have  $\lvert N_{L/\mathbb{Q}}(u)\rvert=1$. The unit  $u$ of  $L$ and its conjugate  $\sigma_{1}(u)$ are units of  $L$, so  $u\sigma_{1}(u)$ is a unit of  $L$. Since  $u\sigma_{1}(u)=N_{L/\mathbb{Q}(\sqrt{d_{1}})}(u)\in\mathbb{Q}(\sqrt{d_{1}})$, it is also a unit of $\mathbb{Q}(\sqrt{d_{1}})$. Hence  $u\sigma_{1}(u)=\pm u_{1}^{m_{1}}$ for some $m_{1}\in\mathbb{Z}$. Similarly, $u\sigma_{2}(u)=\pm u_{2}^{m_{2}}$ and $u\sigma_{1}(u)=\pm u_{3}^{m_{3}}$ for some $m_{2},m_{3}\in\mathbb{Z}$. Then we have
$u^{2}=\lvert u^{2}\cdot N_{L/\mathbb{Q}}(u)\rvert=\lvert u^{3}\sigma_{1}(u)\sigma_{2}(u)\sigma_{3}(u)\rvert=\lvert\pm u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}\rvert=u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}$
and therefore $u^{2}\in E$. ∎
[…]
@@
[…]
where $X_{1}=\log(u_{2})\log(u_{3}),X_{2}=\log(u_{1})\log(u_{3}),X_{3}=\log(u_{1})\log(u_{2})$ and the order of the basis of $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma_{1}},\delta^{\sigma_{2}}\wedge\delta^{\sigma_{3}},\delta^{id}\wedge\delta^{\sigma_{3}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{2}},\delta^{id}\wedge\delta^{\sigma_{2}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{3}}$. Note that $X_{1}>X_{2}>X_{3}>0$.
For $w\in\bigwedge^{2}\operatorname{LOG}(E)$$w$ can be written as
@@
$+\lvert n_{1}X_{1}+n_{3}X_{3}\rvert+\lvert n_{1}X_{1}-n_{3}X_{3}\rvert)$
$=$  $\,4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
@@
$\lVert w\rVert_{1}$ $=4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
@@
(3.4) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 1.697.$
@@
(3.5) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(\frac{3+\sqrt{13}}{2})\approx 2.300$
@@
by (3.3). Excluding $\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $\mathbb{Q}(\sqrt{5},\sqrt{13})$, $\mathbb{Q}(\sqrt{3},\sqrt{5})$ has the smallest possible value of  $\log(u_{1})\log(u_{2})$, which is  $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})$. This is from the Lemma 3.3 below and the fact that $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})<\log(1+\sqrt{2})\log(\frac{3+\sqrt{13}}{2})$. Hence from (3.6), we see that
$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})\approx 1.267.$
@@
$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
@@
For square-free integer $m$, let $v_{m}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{m})$. Then  $\frac{1+\sqrt{5}}{2},1+\sqrt{2},\frac{3+\sqrt{13}}{2},2+\sqrt{3}$ are the four smallest possible values of $v_{m}$.

### v2

@@ Abstract.
The conjecture due to Bertrand and Rodriguez Villegas asserts that the 1-norm of the nonzero element in an exterior power of the units of a number field has a certain lower bound. For the exterior square case of totally real quartic extensions of the rationals, Costa and Friedman gave a lower bound of 0.802. We prove that the bound can be improved to 1.134 when the extension is further assumed to be Galois.
@@
Bertrand–Rodriguez Villegas conjecture, units, heights, Lehmer’s conjecture
@@ 1. Introduction
While Dirichlet’s theorem determines the rank of the unit group in a given algebraic number field in terms of its Archimedean places, less is known about its geometry when viewed as a lattice in a Euclidean space via the associated logarithm embedding. The resulting lattice and its exterior powers are conjecturally subject to certain numerical constraints [1, 2]. We begin by introducing notation to clarify the nature of this numerical constraint.
@@
to get an orthonormal basis  $\{\delta^{I}\}_{I\in\mathcal{A}_{L}^{[j]}}$ of  $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$. For $w=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}\delta^{I}\in\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$, define its 1-norm and 2-norm as
$\lVert w\rVert_{1}:=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}\lvert c_{I}\rvert,\qquad\lVert w\rVert_{2}:=\bigg(\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}^{2}\bigg)^{1/2}$
respectively.
Our aim in this paper is to investigate a lower bound for  $\lVert w\rVert_{1}$ for a nonzero element  $w$ of  $\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. For a positive integer $n$, let $L_{n}$ denote the set of subfields of $\mathbb{R}$ that are Galois extension of $\mathbb{Q}$ of degree $n$. For such fields, $\operatorname{rank}({\mathcal{O}_{L}^{*}})=n-1$ by Dirichlet’s unit theorem. For $1\leq j\leq n-1$, define a real number $A_{n,j}$ as
@@
This constant is directly related to the real Galois extension case of Bertrand’s and Rodriguez Villegas’ conjecture. The conjecture stated in [2] is as follows.
@@
There exist two absolute constants
@@
$A_{4,2}>2\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 0.802,$
as explained in § 2. Throughout this paper, we write $\log^{2}(x)$ to mean $(\log(x))^{2}$. To the best of the authors’ knowledge, this was the largest known lower bound for $A_{4,2}$.
@@
$\lVert w\rVert_{1}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 1.134$
@@
$A_{4,2}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 1.134.$
As demonstrated in [2], the Bertrand–Rodriguez Villegas conjecture is equivalent to Lehmer’s conjecture [7] when $j=1$, and to Zimmert’s theorem on regulators [9] when $j=\operatorname{rank}({\mathcal{O}_{L}^{*}})$. Hence the case of  $n=4,j=2$ is neither covered by Lehmer’s conjecture nor Zimmert’s theorem. While our result only covers this specific case, we obtained a better lower bound, which is $\sqrt{2}$ times the previous known lower bound. In fact, for $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $w=\pm\operatorname{LOG}(\tfrac{1+\sqrt{5}}{2})\wedge\operatorname{LOG}(1+\sqrt{2})$, we have  $\lVert w\rVert_{1}=8\log(\tfrac{1+\sqrt{5}}{2})\log(1+\sqrt{2})$ and thus
$A_{4,2}\leq 8\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 3.393.$
Hence our lower bound is no less than a third of an upper bound for $A_{4,2}$.
The proof of the theorem is done separately for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is a Klein four-group, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is cyclic. For the Klein four-group case, we define a subgroup $E$ generated by the units of the quadratic subfields of $L$. Using the Galois module structure of $E$, we give a lower bound for  $\lVert w\rVert_{1}$ for nonzero element  $w$ of $\bigwedge^{2}\operatorname{LOG}(E)$ and extend this result to $\mathcal{O}_{L}^{*}$}. For the cyclic group case, we use the structure of  $\mathcal{O}_{L}^{*}$} given by Hasse [5] to obtain a basis of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. We then use the result of Pohst [8] to compute the lower bound.
In § 2 we derive the best lower bound for  $A_{4,2}$ known so far, from the work of Costa and Friedman [3]. In the remaining sections, we prove the main theorem to give the better lower bound for $A_{4,2}$. The case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group is covered in § 3, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group is covered in § 4.
@@ Acknowledgement
The work of S.S. was supported by the College of Natural Sciences Undergraduate Internship Program from Seoul National University. D.K. is supported by the National Research Foundation of Korea (NRF)111the grants No. RS-2023-00301976 and No. 2020R1C1C1A0100681913. We also thank the referee for corrections, including those concerning the cyclic case.
@@ 2. The lower bound due to Costa and Friedman
In this section, we derive the largest lower bound for  $A_{4,2}$ known so far, to the best of the authors’ knowledge. Using Pohst’s result [8], Costa and Friedman [3] showed that for independent elements $\epsilon_{1},...,\epsilon_{j}$ of $\mathcal{O}_{L}^{*}$}, the inequality
$\lVert\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{j})\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{j/2}$
holds for $1\leq j<n$, where $n=[L:\mathbb{Q}]$ and  $\gamma_{j}$ is Hermite’s constant in dimension $j$. If $j=n-2=\operatorname{rank}({\mathcal{O}_{L}^{*}})-1$, every element of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ can be written as  $d\cdot\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{n-2})$ where $d\in\mathbb{Z}$ and $\operatorname{LOG}(\epsilon_{1}),...,\operatorname{LOG}(\epsilon_{n-1})$ are the basis of $\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, as shown in Lemma 28 of [2]. Therefore, for any nonzero element  $w$ of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, the inequality
$\lVert w\rVert_{1}\geq\lVert w\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{j/2}$
@@
$A_{n,n-2}>\left(\frac{n}{\gamma_{n-2}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{\frac{n-2}{2}}$
@@
$A_{4,2}>2\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 0.802.$
@@ 3. The Klein four-group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Let $\mathbb{Q}(\sqrt{d_{1}}),\mathbb{Q}(\sqrt{d_{2}}),\mathbb{Q}(\sqrt{d_{3}})$ be three quadratic subfields of $L$, where $d_{1},d_{2},d_{3}$ are square-free integers. Then we can write $L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$. For $i=1,2,3$, let $\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))=\{1,\sigma_{i}\}$. Let  $u_{i}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{d_{i}})$. Without loss of generality, let $1<u_{1}<u_{2}<u_{3}$. If $i$ and $j$ are distinct integers in $\{1,2,3\}$, since $\sqrt{d_{j}}\not\in\mathbb{Q}(\sqrt{d_{i}})$$\sigma_{i}\in\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))$ sends $\sqrt{d_{j}}$ to its conjugate $\sqrt{d_{j}}$. Thus
@@
Let $E=\{u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}|m_{1},m_{2},m_{3}\in\mathbb{Z}\}$. Then we have the following lemmas, which are immediate consequences of Kubota’s classification of fundamental systems of units in real biquadratic fields [6, Satz 1].
@@
We first compute the 1-norm of the elements of $\bigwedge^{2}\operatorname{LOG}(E)$. $\operatorname{LOG}(E)$ is generated by $\operatorname{LOG}(u_{i})$’s. Using (3.1), we can compute $\operatorname{LOG}(u_{i})$’s in terms of $u_{i}$’s as follows:
@@
$\operatorname{LOG}(u_{2})\wedge\operatorname{LOG}(u_{3})=(\ \ \ \ \ 0,\ \ \ \ \ 0,\ \ 2X_{1},\ \ 2X_{1},-2X_{1},-2X_{1})$
$\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{3})=(-2X_{2},-2X_{2},\ \ 2X_{2},-2X_{2},\ \ \ \ \ 0,\ \ \ \ \ 0)$
$\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})=(-2X_{3},\ \ 2X_{3},\ \ \ \ \ 0,\ \ \ \ \ 0,\ \ 2X_{3},-2X_{3})$
where $X_{1}=\log(u_{2})\log(u_{3}),X_{2}=\log(u_{1})\log(u_{3}),X_{3}=\log(u_{1})\log(u_{2})$ and the order of the basis of $\bigwedge^{2}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma_{1}},\delta^{\sigma_{2}}\wedge\delta^{\sigma_{3}},\delta^{id}\wedge\delta^{\sigma_{3}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{2}},\delta^{id}\wedge\delta^{\sigma_{2}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{3}}$. Note that $X_{1}>X_{2}>X_{3}>0$.
For $\in\bigwedge^{2}\operatorname{LOG}(E)ww$ can be written as
@@
[…]
@@
[…]
@@
(3.4) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 1.697.$
@@
(3.5) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log\left(\tfrac{3+\sqrt{13}}{2}\right)\approx 2.300$
@@
by (3.3). Excluding $\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $\mathbb{Q}(\sqrt{5},\sqrt{13})$, $\mathbb{Q}(\sqrt{3},\sqrt{5})$ has the smallest possible value of  $\log(u_{1})\log(u_{2})$, which is  $\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})$. This is from the Lemma 3.3 below and the fact that $\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})<\log(1+\sqrt{2})\log\left(\frac{3+\sqrt{13}}{2}\right)$. Hence from (3.6), we see that
$\lVert w\rVert_{1}\geq 2\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})\approx 1.267.$
@@
$\lVert w\rVert_{1}\geq 2\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})>2\sqrt{6}\log\left(\tfrac{1+\sqrt{5}}{2}\right)^{2}$
@@
For square-free integer $m$, let $v_{m}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{m})$. Then  $\tfrac{1+\sqrt{5}}{2},1+\sqrt{2},\tfrac{3+\sqrt{13}}{2},2+\sqrt{3}$ are the four smallest possible values of $v_{m}$.